# UTG: Towards a Unified View of Snapshot and Event Based Models for Temporal Graphs

### v1

@@ Abstract
Temporal graphs have gained increasing importance due to their ability to model dynamically evolving relationships. These graphs can be represented through either a stream of edge events or a sequence of graph snapshots. Until now, the development of machine learning methods for both types has occurred largely in isolation, resulting in limited experimental comparison and theoretical cross-pollination between the two. In this paper, we introduce Unified Temporal Graph (UTG), a framework that unifies snapshot-based and event-based machine learning models under a single umbrella, enabling models developed for one representation to be applied effectively to datasets of the other. We also propose a novel UTG training procedure to boost the performance of snapshot-based models in the streaming setting. We comprehensively evaluate both snapshot and event-based models across both types of temporal graphs on the temporal link prediction task. Our main findings are threefold: first, when combined with UTG training, snapshot-based models can perform competitively with event-based models such as TGN and GraphMixer even on event datasets. Second, snapshot-based models are at least an order of magnitude faster than most event-based models during inference. Third, while event-based methods such as NAT and DyGFormer outperforms snapshot-based methods on both types of temporal graphs, this is because they leverage joint neighborhood structural features thus emphasizing the potential to incorporate these features into snapshot-based models as well. These findings highlight the importance of comparing model architectures independent of the data format and suggest the potential of combining the efficiency of snapshot-based models with the performance of event-based models in the future.
@@ 1 Introduction
Distinct Evaluation Settings. Another obstacle to comparing snapshot and event-based methods is their distinct evaluation settings. Snapshot-based methods have been primarily tested under the deployed setting [30, 15], where the test set information is strictly not available to the model, and training set information is used for prediction. In contrast, event-based models are designed for the streaming setting [18, 7], where streaming predictions allow the model to use recently observed information, enabling event-based models to update their node representations at test time.
@@
Reproducibility: The datasets used in this work are public. The code for reproducing the experiments are available in the supplementary materials.
@@ 2 Related Work
Recently, many ML methods were developed for temporal graphs. The well-adopted categorization by Kazemi et al. [8] considers two types of temporal graphs: Discrete Time Dynamic Graphs (DTDGs) and Continuous Time Dynamic Graphs (CTDGs). More detailed discussions on related work can be found in Appendix F.
@@
Discrete Time Dynamic Graphs. Early methods often represent temporal graphs as a sequence of graph snapshots while adapting common graph neural networks such as Graph Covolution Network (GCN) [9] used in static graphs for DTDGs. For example, EGCN [15] employs a Recurrent Neural Network (RNN) to evolve the parameters of a GCN over time. In comparison, GCLSTM [2] learns the graph structure via a GCN while capturing temporal dependencies with an LSTM network [5]. Pytorch Geometric-Temporal (PyG Temporal) [20] is a comprehensive framework that facilitate neural spatiotemporal signal processing which implements existing work such as EGCN and GCLSTM in an efficient manner. HTGN [30] utilizes hyperbolic geometry to better capture the complex and hierarchical nature of the evolving networks. Recently, You et al. [32] introduced a novel live-update setting where GNNs are always trained on the most recent observed snapshot after making predictions. In comparison, the streaming setting in this work allows the model to use observed snapshots for forward pass but no training are permitted on test set.
@@
Continuous Time Dynamic Graphs. Event-based methods process temporal graphs as a stream of timestamped edges. DyRep [23] and JODIE [10] are two pioneering work on CTDGs. TGAT [28] is one of the first work for studying inductive representation learning on temporal graphs. Rossi et al. [18] introduce Temporal Graph Networks (TGNs), a generic inductive framework of Temporal Graph Networks, showing DyRep, JODIE and TGAT as its special cases. Methods such as CAWN [26] and NAT [12] both focuses on learning the joint neighborhood of the two nodes of interest in the link prediction task. CAWN focuses on learning from temporal random walks while NAT is a neighborhood-aware temporal network model that introduces a dictionary-type neighborhood representation for each node. TCL [25] and DyGFormer [33] applies transformer based architecture on CTDGs, inspired by the success of transformer based architectures on time series [27], images [4] and natural language processing [1].
Evaluations in Graph Machine Learning. Datasets spanning a multitude of domains could have distinct underlying mechanisms of link formation; thus, it is important to explore link prediction across diverse datasets [13]. The Open Graph Benchmark [6] provides a large collection of datasets and tasks for benchmarking on static graphs. Recently, Huang et al. [7] proposed the Temporal Graph Benchmark (TGB), a collection of large-scale and realistic datasets from distinct domains for both link and node level tasks on CTDGs. In this work, we experiment on the tgbl-wiki and tgbl-review dataset to compare both snapshot and event-based methods. Poursafaei et al. [16] suggested that the performance of temporal link prediction is often over-optimistic due to the easy negative samples used. Thus, they propose to use the historical negatives which are past edges absent in the current step for evaluation. In addition, a simple heuristic baseline called Edgebank is proposed as a strong contender with fast computational time. In this work, we also incorporate historical negative samples for our evaluation and adding EdgeBank in the comparison.
@@ 3 Preliminaries
Without loss of generality, one can normalize the timestamps in $\mathcal{G}$ from $[t_{0},t_{k}]$ to $[0,1]$ by applying $t=\frac{t-t_{0}}{t_{k}}\forall t\in[t_{0},t_{k}]$. Real world temporal networks can be broadly classified into two inherent types based on the nature of their edges: spontaneous networks and relationship networks. Examples of spontaneous networks include transaction networks, retweet networks, Reddit networks, and other activity graphs. Here, the edges are spontaneous thus resulting in the start time and end time of an edge being the same, i.e. $t^{s}=t^{e}$. This formulation is inline with related studies in [16, 8, 7, 21]. For relationship networks such as friendship networks, contact networks, and collaboration networks, the edges often persist over a period of time resulting in $t^{s}\neq t^{e}$.
@@ 4 UTG Framework
Figure 1: Illustration of the UTG framework. The input graph is processed by the UTG input mapper to generate the appropriate input data format for event-based or snapshot-based models. The model predictions are then processed by the UTG output mapper for the respective prediction task.
In this section, we present the Unified Temporal Graph (UTG) framework which aims to unify snapshot-based and event-based temporal graph models under the same framework, enabling temporal graph models to be applied to both CTDGs and DTDGs. UTG has two key components: input mapper and output mapper. Input mapping converts the input temporal graph into the appropriate representation needed for a given method, i.e. snapshots or events. Output mapper transforms the prediction of the model to the required time granularity of the task. Figure 1 shows the workflow of UTG framework. UTG enables any temporal graph learning methods to be applied to any input temporal graph via UTG input mapper and UTG output mapper.
@@ 4.1 UTG Input Mapper
Both snapshot and event-based TG methods require specific input data format. For snapshot-based models, discretizing CTDG data into a sequence of snapshots is required (Section 4.1.1). For event-based models, DTDG snapshots need to be converted into batches of events (Section 4.1.2).
4.1.1 Converting CTDG to Snapshots
@@
where $\mathbf{G}_{\tau_{i}}^{\tau_{j}}$ is defined as the aggregated graph snapshot containing all edges that have a start time $t_{s}\leq\tau_{j}$ and an end time $t_{e}\geq\tau_{i}$, i.e. edges that are present solely within the $[{\tau_{i}},{\tau_{j}}]$ interval.
@@
4.1.2 Converting DTDG to Events
While it may seem straightforward to convert DTDG to events — simply create one event with timestep $t$ for each edge in snapshot $G_{t}$ — some subtleties related to batch training and memory update of event-based models have to be considered to avoid data leakage. Event-based models often receive batches of events (or edges) with a fixed dimension as inputs [18, 33, 12]. In discrete-time dynamic graphs, all edges in a snapshot have the same timestamp and are assumed to arrive simultaneously. Therefore, using a fixed batch size can result in splitting the snapshot into multiple batches. Because models in the streaming setting [7] such as TGN [18] and NAT [12] update their representation of the temporal graph at the end of each batch, predicting a snapshot across multiple batches leads to data leakage: a portion of the edges from the snapshot is used to predict other (simultaneous) edges from the same snapshot. To avoid data leakage on DTDGs, we ensure that each snapshot is contained in a single batch for event-based models [^1] In case the resulting batch would not fit in memory, one can delay the memory update (and parameter updates during training) only after all edges from the current snapshot have been processed, something akin to gradient accumulation..
@@ 4.2 UTG Output Mapper
By applying zero-order hold for snapshot-based models, the predictions can now be broadcasted for a period of time (specifically for the duration of a given snapshot $[\tau_{i},\tau_{j}]$). Therefore, it is now possible to utilize snapshot-based models on continuous-time dynamic graphs.
5 Improving Snapshot-based Models
Here, we discuss the changes to the snapshot-based methods in the UTG framework.
5.1 Streaming Setting
@@
Figure 2: UTG training and evaluation workflow. UTG training enables snapshot-shot based models to incorporate information from observed snapshots at inference time.
@@
Event-based models often evaluate with the streaming setting [18, 33, 26, 12]. In this setting, information from the previously observed batches of events (or graph snapshot) can be used to update the model however no information from the test set is used to train the model.
@@
We also adopt the streaming setting which closely resembles real-world settings where after the model is trained, it is required to incorporate newly observed information into its predictions. Figure 2 shows the evaluation pipeline used for snapshot-based models in UTG. After a snapshot is observed, it can be used to update the node representation of the snapshot-based models for the prediction of the next snapshot (only forward pass for inference). More discussion on evaluation settings are shown in Appendix E.
5.2 UTG Snapshot Training
Figure 2 illustrates the UTG training procedure for snapshot-based models. In standard training, e.g. with Pytorch Geometric Temporal [19], the model uses snapshots up to $\mathbf{G}_{t}$ to predict edges for the same time $t$, leading to information leakage as the model predicts edges that were part of its input. Moreover, the gradients from all training snapshots are accumulated before performing backpropagation to update the model weights. In contrast, UTG training introduces two key differences: (1) only snapshots up to $\mathbf{G}_{t-1}$ are used to predict the edges for time $t$, and (2) it performs backpropagation at each snapshot rather than accumulating gradients for all training snapshots. In Section 6.4, we conduct an ablation study on the GCLSTM model to demonstrate the benefits of UTG training.
@@ 6 Experiments
Datasets. In this work, we consider five discrete-time dynamic graph datasets and three continuous-time dynamic graph datasets. tgbl-wiki and tgbl-review are datasets from TGB [7] while the rest are found in Poursafaei et al. [16]. The dataset statistics are shown in Table 1. The time granularity or discretization level of each DTDG dataset is selected as the finest time granularities where there are no time gaps (see Section 4.1.1). The surprise index is defined as $surprise=\frac{|E_{\text{test}}\setminus E_{\text{train}}|}{E_{\text{test}}}$ [16] which measures the proportion of unseen edges in the test set when compared to the training set.
@@
Evaluation Setting. For each dataset, we generate a fixed set of negative samples for each positive edge consisting of 50% historical negatives and 50% random negative, same as in TGB [7]. For DTDG datasets, we generate 1000 negative samples per positive edge. For TGB datasets, we use the same set of negatives provided in TGB and for Reddit, we generate 1000 negatives similar to before. For graphs with less than 1k nodes, we generate negative samples equal to number of nodes. We follow the streaming setting where the models are allowed to update their representation at test time while gradient updates are not permitted. For DTDG datasets, we select the best results from learning rate $0.001$ or $0.0002$. For CTDG datasets, we report the results from TGB [7] where available or by learning rate $0.0002$.
Compared Methods. We compare four event-based methods including TGN [18], DyGFormer [33], NAT [12] and GraphMixer [3]. We also include Edgebank [16], a scalable and non-parameteric heuristics. In addition, we compare three existing snapshot-based methods including HTGN [30], GCLSTM [2] and EGCNo [15]. Lastly, we adapt a common 2-layer (static) GCN [9] under the UTG framework to demonstrate the flexibility of UTG (without a recurrent module). If a method runs out of memory on a NVIDIA A100 GPU (40GB memory), it is reported as out of memory (OOM). If a method runs for more than 5 days, it is reported as out of time (OOT)
6.1 Comparing Event-Based with Snapshot-Based Models
@@
Table 2: Test MRR comparison for snapshot and event-based methods on DTDG datasets, results reported from 5 runs. Top three models are coloured by First, Second, Third.
@@
l | l | l l l l l
Method UCI Enron Contacts Social Evo. MOOC
@@
DTDG Results. Table 6.1 shows the performance of all methods on the DTDG datasets. Surprisingly, we find that event-based methods achieve state-of-the-art performance on the DTDG datasets, particularly with DyGFormer and NAT consistently outperforming other methods.
With the improvements from UTG, snapshot-based models can obtain competitive performance on datasets such as Enron and Social Evo. Interestingly, even the simple GCN with UTG training can achieve second place performance on the Social Evo. dataset. Note that this dataset has the lowest surprise out of all datasets meaning the majority of test set edges have been observed during training, possibly explaining the strong performance of GCN in this case. Lastly, on the MOOC dataset which has the largest number of nodes out of all DTDG datasets, both DyGformer and GraphMixer ran out of memory (OOM) showing their difficulty in scaling with the number of nodes in a snapshot.
Table 3: Test MRR comparison for snapshot and event-based methods on CTDG datasets, results reported from 5 runs. Top three models are coloured by First, Second, Third.
@@
CTDG Results. Table 6.1 shows the performance of all methods on the CTDG datasets. Similar to DTDG datasets, DyGformer and NAT retains competitive performance here. On the tgbl-wiki and Reddit dataset, HTGN, a snapshot-based model is able to outperform widely-used TGN and GraphMixer model. This shows that learning from the discretized snapshots can be effective even on CTDG datasets. However, snapshot-based models have lower performance on the tgbl-review dataset where the surprise index is high. This shows that the inductive reasoning capability on snapshot-based models should be further improved to generalize to unseen edges.
6.2 Computational Time Comparison
@@
Figure 3: Snapshot-based models are at least an order of magnitude faster than event-based models for inference.
@@
Figure 2(a) and Figure 2(b) shows the test inference time comparison for all methods on the Social Evo. and the tgbl-review dataset respectively. The test time for each dataset is reported in Appendix D. Overall, we observe that snapshot-based methods are at least an order of magnitude faster than most event-based methods. In comparison, high performing model such as DyGformer has significantly higher computational time thus limiting its scalability to large datasets. One important future direction is to combine the predictive power of event-based models such as NAT and DyGFormer with that of the efficiency of the snapshot-based model for improved scalability.
6.3 Discussion
Event-based methods such as NAT and DyGFormer tend to perform best on both CTDG and DTDG, potentially leading to the premature conclusion that event-based modeling is the preferred paradigm and that snapshot-based models should be avoided. However, the superior performance of NAT and DyGFormer could be primarily due to their ability to leverage joint neighborhood structural features, specifically the common neighbors between the source and destination nodes of a link [12, 33]. This approach has been shown to be fundamental for achieving competitive link prediction on static graphs [34, 35]. In contrast, none of the existing snapshot-based methods incorporate joint neighborhood structural features. Therefore, the performance difference could be mainly attributed to this factor rather than an intrinsic difference between event-based and snapshot-based models. This is confirmed by the fact that event-based models such as TGN and GraphMixer, which omits these features, have no clear performance advantage over snapshot-based methods. Moreover, snapshot-based methods are more computationally efficient and might be preferred when efficiency is important. These considerations suggest that both event-based and snapshot-based methods have their own merit. We believe that combining the strengths of both approaches will be an important direction for future research.
6.4 UTG Training Ablation Study

### v2

@@ Abstract
Many real world graphs are inherently dynamic, constantly evolving with node and edge additions. These graphs can be represented by temporal graphs, either through a stream of edge events or a sequence of graph snapshots. Until now, the development of machine learning methods for both types has occurred largely in isolation, resulting in limited experimental comparison and theoretical cross-pollination between the two. In this paper, we introduce Unified Temporal Graph (UTG), a framework that unifies snapshot-based and event-based machine learning models under a single umbrella, enabling models developed for one representation to be applied effectively to datasets of the other. We also propose a novel UTG training procedure to boost the performance of snapshot-based models in the streaming setting. We comprehensively evaluate both snapshot and event-based models across both types of temporal graphs on the temporal link prediction task. Our main findings are threefold: first, when combined with UTG training, snapshot-based models can perform competitively with event-based models such as TGN and GraphMixer even on event datasets. Second, snapshot-based models are at least an order of magnitude faster than most event-based models during inference. Third, while event-based methods such as NAT and DyGFormer outperforms snapshot-based methods on both types of temporal graphs, this is because they leverage joint neighborhood structural features thus emphasizing the potential to incorporate these features into snapshot-based models as well. These findings highlight the importance of comparing model architectures independent of the data format and suggest the potential of combining the efficiency of snapshot-based models with the performance of event-based models in the future.
@@ 1 Introduction
Distinct Evaluation Settings. Another obstacle to comparing snapshot and event-based methods is their distinct evaluation settings. Snapshot-based methods have been primarily tested under the deployed setting [12, 14]  [^1] An exception to this is ROLAND [18], which proposed the live-update setting (see Appendix B for a more detailed discussion)., where the test set information is strictly not available to the model, and training set information is used for prediction. In contrast, event-based models are designed for the streaming setting [15, 19], where streaming predictions allow the model to use recently observed information, enabling event-based models to update their node representations at test time.
@@
Reproducibility: The code and data for this project is publicly available on Github:
https://github.com/shenyangHuang/UTG
@@ 2 Related Work
Holme et al. [20] provided a general overview of the many types of real world temporal networks showing that temporal graphs are ubiquitous in many applications. Recently, many ML methods were developed for temporal graphs. The well-adopted categorization by Kazemi et al. [10] are defined by the two types of temporal graphs: Discrete Time Dynamic Graphs (DTDGs) and Continuous Time Dynamic Graphs (CTDGs). Due to difference in input data format, empirical comparison between methods designed for CTDGs and DTDGs are under-explored and these two categories are often considered distinct lines of research despite many similarities in models’ design. More detailed discussions on related work can be found in Appendix C.
@@
Discrete Time Dynamic Graphs. Early methods often represent temporal graphs as a sequence of graph snapshots while adapting common graph neural networks such as Graph Covolution Network (GCN) [1] used in static graphs for DTDGs. For example, EGCN [14] employs a Recurrent Neural Network (RNN) to evolve the parameters of a GCN over time. In comparison, GCLSTM [13] learns the graph structure via a GCN while capturing temporal dependencies with an LSTM network [21]. Pytorch Geometric-Temporal (PyG Temporal) [22] is a comprehensive framework that facilitate neural spatiotemporal signal processing which implements existing work such as EGCN and GCLSTM in an efficient manner. HTGN [12] utilizes hyperbolic geometry to better capture the complex and hierarchical nature of the evolving networks. Recently, You et al. [18] introduced a novel live-update setting where GNNs are always trained on the most recent observed snapshot after making predictions. In comparison, the streaming setting in this work allows the model to use observed snapshots for forward pass but no training are permitted on test set. Zhu et al. [23] designed the WinGNN framework for the live-update setting where a simple GNN with meta-learning strategy is used in combination with a novel random gradient aggregation scheme, removing the need for temporal encoders.
@@
Continuous Time Dynamic Graphs. Event-based methods process temporal graphs as a stream of timestamped edges. DyRep [24] and JODIE [9] are two pioneering work on CTDGs. TGAT [25] is one of the first work for studying inductive representation learning on temporal graphs. Rossi et al. [15] introduce Temporal Graph Networks (TGNs), a generic inductive framework of Temporal Graph Networks, showing DyRep, JODIE and TGAT as its special cases. Methods such as CAWN [26] and NAT [17] both focuses on learning the joint neighborhood of the two nodes of interest in the link prediction task. CAWN focuses on learning from temporal random walks while NAT is a neighborhood-aware temporal network model that introduces a dictionary-type neighborhood representation for each node. TCL [27] and DyGFormer [28] applies transformer based architecture on CTDGs, inspired by the success of transformer based architectures on time series [29], images [30] and natural language processing [31]. Despite the promising performance of event-based methods, recent work showed significant limitations in the standard link prediction evaluation due to the simplicity of negative samples used for evaluation [16]. To improve the evaluation for CTDG, Huang et al. [19] proposed the Temporal Graph Benchmark (TGB), a collection of large-scale and realistic datasets from distinct domains for both link and node level tasks.
@@ 3 Preliminaries
Without loss of generality, one can normalize the timestamps in $\mathcal{G}$ from $[t_{0},t_{k}]$ to $[0,1]$ by applying $t=\frac{t-t_{0}}{t_{k}}\forall t\in[t_{0},t_{k}]$. Real world temporal networks can be broadly classified into two inherent types based on the nature of their edges: transient networks and persistent networks. Examples of spontaneous networks include transaction networks, retweet networks, Reddit networks, and other activity graphs. Here, the edges are spontaneous thus resulting in the start time and end time of an edge being the same, i.e. $t^{s}=t^{e}$. This formulation is inline with related studies in [16, 10, 19, 32]. For relationship networks such as friendship networks, contact networks, and collaboration networks, the edges often persist over a period of time resulting in $t^{s}\neq t^{e}$.
@@ 4 UTG Framework
Figure 1: Illustration of the UTG framework. The input graph is processed by the UTG input mapper to generate the appropriate input data format for TG models. The model predictions are then processed by the UTG output mapper for prediction.
In this section, we present the Unified Temporal Graph (UTG) framework which aims to unify snapshot-based and event-based temporal graph models under the same framework, enabling temporal graph models to be applied to both CTDGs and DTDGs. UTG has two key components: input mapper and output mapper. Input mapping converts the input temporal graph into the appropriate representation needed for a given method, i.e. snapshots or events. Output mapper transforms the prediction of the model to the required time granularity of the task. Figure 1 shows the workflow of UTG framework. UTG enables any temporal graph learning methods to be applied to any input temporal graph.
@@ 4.1 UTG Input Mapper
Both snapshot and event-based TG methods require specific input data format. For snapshot-based models, discretizing CTDG data into a sequence of snapshots is required. For event-based models, DTDG snapshots need to be converted into batches of events.
Converting CTDG to Snapshots
@@
where $\mathbf{G}_{\tau_{i}}^{\tau_{j}}$ is defined as the aggregated graph snapshot containing all edges that have a start time $t_{s}<\tau_{j}$ and an end time $t_{e}\geq\tau_{i}$, i.e. edges that are present solely within the $[{\tau_{i}},{\tau_{j}})$ interval.
@@
Converting DTDG to Events
While it may seem straightforward to convert DTDG to events — simply create one event with timestep $t$ for each edge in snapshot $G_{t}$ — some subtleties related to batch training and memory update of event-based models have to be considered to avoid data leakage. Event-based models often receive batches of events (or edges) with a fixed dimension as inputs [15, 28, 17]. In discrete-time dynamic graphs, all edges in a snapshot have the same timestamp and are assumed to arrive simultaneously. Therefore, using a fixed batch size can result in splitting the snapshot into multiple batches. Because models in the streaming setting [19] such as TGN [15] and NAT [17] update their representation of the temporal graph at the end of each batch, predicting a snapshot across multiple batches leads to data leakage: a portion of the edges from the snapshot is used to predict other (simultaneous) edges from the same snapshot. To avoid data leakage on DTDGs, we ensure that each snapshot is contained in a single batch for event-based models [^2] In case the resulting batch would not fit in memory, one can delay the memory update (and parameter updates during training) only after all edges from the current snapshot have been processed, something akin to gradient accumulation.. Note that the issue of splitting edges that share the same timestamp into multiple batches can exist in general CTDG datasets as well, more likely for datasets with a large burst of edges at a single timestamp.
@@ 4.2 UTG Output Mapper
By applying zero-order hold for snapshot-based models, the predictions can now be broadcasted for a period of time (specifically for the duration of a given snapshot $[\tau_{i},\tau_{j}]$). Therefore, it is now possible to utilize snapshot-based models on continuous-time dynamic graphs. Note that often snapshot-based model are designed to predict for the immediate next snapshot and not capable of predicting snapshots in more distant future. Therefore, inherently their ability to predict events in the far future is limited when compared to event-based model that explicitly takes a timestamp as input. With zero-order hold, we assume that the event to predict next is within the next snapshot to circumvallate the aforementioned limitation of snapshot-based model. To achieve this, we select the finest time granularity on the CTDG datasets which results in no time gap to construct snapshots.
4.3 Streaming Evaluation
@@
Figure 2: UTG training and evaluation workflow. UTG training enables snapshot-based models to perform better for the streaming setting.
@@
Event-based models often evaluate with the streaming setting [15, 28, 26, 17]. In this setting, information from the previously observed batches of events (or graph snapshot) can be used to update the model however no information from the test set is used to train the model. In comparison, snapshot-based models are often evaluated in either the live-update setting [18] or the deployed setting [14, 12]. Detailed differences between evaluation settings are discussed in Appendix B.
@@
In this work, to provide a unified comparison, we focus on the widely used streaming setting for experimental evaluation as it closely resembles real-world settings where after the model is trained, it is required to incorporate newly observed information into its predictions. Note that the UTG framework can also be applied to test under the deployed setting with little changes. For the live-update setting, the changes are required for each model’s training procedure as the training set is not split chronologically but rather being a subgraph of each observed snapshot. Figure 2 shows the evaluation pipeline used for snapshot-based models in UTG. After a snapshot is observed, it can be used to update the node representation of the snapshot-based models for the prediction of the next snapshot (only forward pass for inference).
4.4 UTG Training for Snapshot-based Models
Here, we discuss the changes to the snapshot-based methods in the UTG framework. Figure 2 illustrates the work flow of UTG training for snapshot-based models. The standard training for snapshot-based models, e.g. with Pytorch Geometric Temporal [33], are designed for tasks such as graph regression or node classification, for the deployed setting. Therefore, the training procedure needs to be adapted accordingly for the link prediction with the streaming setting.
Difference of UTG Training. The first required change is that in standard training, the snapshots up until time $t$ is used as input for predictions at time $t$. This is feasible for node classification and graph regression tasks: the graph structure at time $t$ is used to predict the unknown target labels. However, this is problematic for the link prediction task as the target itself (the graph structure) is not available as input to the model. Therefore, to account for this, we modified the training to use only snapshots up to $\mathbf{G}_{t-1}$ as input to predict the graph structure at the current step $\mathbf{G}_{t}$.
The second change is that, in UTG, the loss is backpropagated to the model at each snapshot. This is in contrast with accumulating the loss through the whole sequence and only backpropagation once at the end of training (as seen in standard training). This change is motivated by the training procedure often seen in event-based models for the streaming setting. For example, in TGN [15], loss on each batch is computed based on information from the previous batch. UTG training enhances the performance of snapshot-based models for the streaming setting and in Section 5, we demonstrate the performance advantage of UTG training on a number of snapshot-based models.
Connection to Truncated Backprop Through Time. By considering the temporal graph as a sequence (of snapshots), the link prediction problem can be seen as a time series prediction task where the information until time $(t-1)$ is used to predict for time $t$. Many snapshot-based models utilize a RNN to model temporal dependency. In this view, accumulating gradients throughout the whole sequence before backpropagation is equivalent to the classical backpropagation through time algorithm [34]. From this perspective, backpropagating the loss at each snapshot in UTG training can be interpreted as using the Truncated Backpropagation Through Time (TBTT) algorithm to train the model, with a window size of one. It is known that TBTT helps circumvent common issues of training RNNs such as exploding memory usage and vanishing gradient problem [35, 36]. Therefore, UTG training might help alleviate the vanishing gradient problem.
@@ 5 Experiments
Datasets. In this work, we consider five discrete-time dynamic graph datasets and three continuous-time dynamic graph datasets. tgbl-wiki and tgbl-review are datasets from TGB [19] while the rest are found in Poursafaei et al. [16]. The dataset statistics are shown in Table 1. The time granularity or discretization level of each DTDG dataset is selected as the finest time granularities where there are no time gaps. The surprise index is defined as $surprise=\frac{|E_{\text{test}}\setminus E_{\text{train}}|}{E_{\text{test}}}$ [16] which measures the proportion of unseen edges in the test set when compared to the training set.
@@
Evaluation Setting. A common approach for evaluating dynamic link prediction tasks is similar to binary classification, where one negative edge is randomly sampled for each positive edge in the test set, and performance is measured using metrics like the Area Under the Receiver Operating Characteristic curve (AUROC) or Average Precision (AP) [15, 25]. However, recent studies have shown that such evaluation is overly simplistic and tends to inflate performance metrics for most models [16, 19]. One main reason is that randomly sampled negative edges are too easy and the more challenging historical negatives (past edges absent in the current timestamp) are rarely sampled [16]. To address these issues, several improvements have been proposed, framing the problem as a ranking task where the model must identify the most probable edge from a large pool of negative samples as well as adding challenging negatives. Therefore, we adopt the improved evaluation methodology used in TGB [19], where link prediction is treated as a ranking problem and the Mean Reciprocal Rank (MRR) metric is applied. This metric calculates the reciprocal rank of the true destination node among a large number of possible destinations.
For each dataset, we generate a fixed set of negative samples for each positive edge consisting of 50% historical negatives and 50% random negative, same as in [19]. For DTDG datasets, we generate 1000 negative samples per positive edge. For TGB datasets, we use the same set of negatives provided in TGB and for Reddit, we generate 1000 negatives similar to before. For graphs with less than 1k nodes, we generate negative samples equal to number of nodes. We follow the streaming setting where the models are allowed to update their representation at test time while gradient updates are not permitted. For DTDG datasets, we select the best results from learning rate $0.001$ or $0.0002$. For CTDG datasets, we report the results from TGB [19] where available or by learning rate $0.0002$.
Compared Methods. We compare four event-based methods including TGN [15], DyGFormer [28], NAT [17] and GraphMixer [37]. We also include Edgebank [16], a scalable and non-parameteric heuristics. In addition, we compare three existing snapshot-based methods including HTGN [12], GCLSTM [13], EGCNo [14], and ROLAND-GRU [18]. Lastly, we adapt a common 2-layer (static) GCN [1] under the UTG framework to demonstrate the flexibility of UTG (without a recurrent module). If a method runs out of memory on a NVIDIA A100 GPU (40GB memory), it is reported as out of memory (OOM). If a method runs for more than 5 days, it is reported as out of time (OOT).
@@
Table 2: Test MRR comparison for snapshot and event-based methods on DTDG datasets, results reported from 5 runs. Top three models are marked by First, Second, Third.
@@
l | l | l l l l l Method UCI Enron Contacts Social Evo. MOOC
@@
ROLAND (UTG) [18] 0.103 $\pm$ 0.011 0.243 $\pm$ 0.017 0.145 $\pm$ 0.002 0.240 $\pm$ 0.005 0.121 $\pm$ 0.003
DTDG Results. Table 5 shows the performance of all methods on the DTDG datasets. Surprisingly, we find that event-based methods achieve state-of-the-art performance on the DTDG datasets, particularly with DyGFormer and NAT consistently outperforming other methods. With the improvements from UTG, snapshot-based models can obtain competitive performance on datasets such as Enron and Social Evo. Interestingly, even the simple GCN with UTG training can achieve second place performance on the Social Evo. dataset. Note that this dataset has the lowest surprise out of all datasets meaning the majority of test set edges have been observed during training, possibly explaining the strong performance of GCN in this case. Lastly, on the MOOC dataset which has the largest number of nodes out of all DTDG datasets, both DyGformer and GraphMixer ran out of memory (OOM) showing their difficulty in scaling with the number of nodes in a snapshot.
Table 3: Test MRR comparison for snapshot and event-based methods on CTDG datasets, results reported from 5 runs. Top three models are marked by First, Second, Third.
@@
ROLAND (UTG) [18] 0.289 $\pm$ 0.003 0.297 $\pm$ 0.006 0.211 $\pm$ 0.006
CTDG Results. Table 5 shows the performance of all methods on the CTDG datasets. Similar to DTDG datasets, DyGformer and NAT retains competitive performance here. On the tgbl-wiki and Reddit dataset, HTGN, a snapshot-based model is able to outperform widely-used TGN and GraphMixer model. This shows that learning from the discretized snapshots can be effective even on CTDG datasets. However, snapshot-based models have lower performance on the tgbl-review dataset where the surprise index is high. This shows that the inductive reasoning capability on snapshot-based models should be further improved to generalize to unseen edges.
@@
Figure 3: Snapshot-based models with UTG training are at least an order of magnitude faster than event-based models for inference.
@@
Computational Time Comparison Figure 2(a) and Figure 2(b) shows the test inference time comparison for all methods on the Social Evo. and the tgbl-review dataset respectively. The test time for each dataset is reported in Appendix A. Overall, we observe that snapshot-based methods are at least an order of magnitude faster than most event-based methods. In comparison, high performing model such as DyGformer has significantly higher computational time thus limiting its scalability to large datasets. One promising direction is to combine the performance of event-based models such as NAT and DyGFormer with that of the efficiency of the snapshot-based models for improved scalability.